# Disentangled Clothed Avatar Generation from Text Descriptions

### v1

@@ 5.1 Generated Avatars & Clothes
Figure 6: Our method is able to generate avatars with varied ethnicities, genders, and clothes.
As demonstrated in Fig. 6, our method generates diverse avatars and various types of clothes with detailed texture. Besides, since the clothes were generated on a base human body, they fit with the intended human naturally.
@@ 5.2 Quality Comparisons
We conducted comparisons on three aspects, static clothed-avatar quality, clothes quality, and animation quality. For the generation quality of clothed avatars, we compared with state-of-the-art text-to-3D avatar generation methods TEXture [66], AvatarCLIP [32] and TADA! [47]. As for clothes generation, we conducted comparisons with two of the best existing text-to-3D methods ProlificDreamer [82] and TextMesh [75]. We then compared animation results in the physical simulation environment [2] with the animatable meshes generated by TEXture [66] and TADA! [47].
@@
Figure 7: Quality comparison of the generated clothed avatars from TEXture [66], AvatarCLIP [32], TADA! [47] and our method with the same prompts.
A visual comparison of the generated avatars is presented in Fig. 7. TEXTure [66] takes a fixed SMPL-X mesh as input and suffers from texture inconsistency under different views. AvatarCLIP [32] often produces geometry artifacts and poor facial details. TADA! [47] tends to generate unnatural geometric artifacts and unaligned outfits, e.g. a hull on the African-American woman’s belly and the white teenage girl’s yellow pants that contradicts the “beige legging” described by the prompt. In comparison, our method is able to produce not only the most realistic human with detailed textures but also high-quality clothes that are semantically aligned with texts.
Clothes Quality. We also evaluate the quality of the separated clothes. We compared with state-of-the-art text-to-3D generation methods ProlificDreamer [82] and TextMesh [75] for standalone clothes generation. As demonstrated in Fig. 8, neither TextMesh [75] or ProlificDreamer [82] could generate meaningful clothes shapes & textures. Besides, they also could not control the position and the size of the generated clothes, and thus could not be fitted to an avatar. Meanwhile, our method is able to produce clothes fitting to a specific avatar ana also with high-quality textures and shapes.
Figure 8: Comparisons of clothes generation quality using the same prompts.
@@
5.3 User Study
A comprehensive user study is conducted to evaluate the generation quality of our pipeline. We designed a questionnaire and presented 16 full-body animation made by TEXture [66], TADA! [47], and our method. The survey was conducted among 26 researchers and artists from either academia or industry. The participants are asked to pick the one with the best semantic consistency, static visual quality, and animation quality. As shown in Tab 1, our model was the most preferred one in all aspects of generation quality.
@@ 5.3 User Study
Table 1: User study from 26 users in terms of visual quality, animation quality, and prompt alignment preference. The number means the percentage of users selecting a specific method.
@@ 5.4 Ablation Study
We evaluated the effects of our albedo smoothness constraints and 2D-pose ControlNet guidance. In Fig. 9, we compared the rendered albedo and RGB images w/ or w/o our albedo smoothness constraint $\mathcal{L}_{a}$. As can be seen from the visual results, the albedo generated without $\mathcal{L}_{a}$ inherent shadows and wrinkles, which would lead to unsatisfactory visual artifacts in animations. In comparison, using $\mathcal{L}_{a}$ largely alleviates the problem in the albedo while still enabling the correct rendering of the shadows and wrinkles on RGB images.
Figure 9: Ablation study on our shading model. For each garment, we showcase the albedo map (left) and the rendered RGB image (right). Our full shading model with albedo smoothness constraint $\mathcal{L}_{a}$ enables the rendering of shadows in wrinkles while alleviating baked-in artifacts of shadows in the albedo.
In Fig. 10, we validated the effectiveness of Openpose-based ControlNet as guidance during our human body generation. As could be observed from the results, generated avatars without ControlNet have oversmoothed textures and geometry. In contrast, our utilization of ControlNet produces significantly finer details in both geometry and texture.
Figure 10: Ablation study on ControlNet as guidance. We compare the avatars generated by the ControlNet with the ones generated by the vanilla Stable diffusion model. The results showed that the ControlNet model with 2D pose conditions generates avatars with more detailed geometry and texture.
@@ 5.5 Applications
Virtual Try-on. Our disentangled human-clothes representation inherently allows us to change the outfit of a certain avatar, or put the same clothes on different avatars.
@@
As depicted in Fig. 11, while previous methods like TADA! [47] also enable virtual try-on, they struggle to maintain consistency in clothing and the human identity during the try-on process. As can be seen from the left half of Fig. 11, when changing clothes for the white teenage boy with TADA!, the facial details and body shape are changed; on the other hand, the colors and patterns of the clothes also changed when putting the same clothes on different avatars with TADA!. On the contrary, our disentangled representation supports fixing the human identity and clothes details when switching outfits.
Clothes Shape Editing Our offset-based model also allows the same outfit design to be fitted on people of different shapes. Under our SO-SMPL representation, we could simply combine the human body mesh parameters of the target avatar: the SMPL-X shape $\beta$ and offset $\mathbf{O}_{\text{h}}$, with the geometry offset $\mathbf{O}_{\text{c}}$ and texture MLP $\Gamma_{\text{c}}$ of the intended clothes. As shown in Fig. 12, clothes generated by our methods could be changed according to different human shapes and sizes without re-generating. This flexibility helps our generated clothes to be re-used on different avatars while maintaining fit with their body types.
Figure 12: Shape editing on clothes. Our SO-SMPL representation allows us to fit generated clothes onto different human shapes while maintaining the original patterns and details.
@@ 6 Limitations&Conlusion
Limitations. Despite promising results, our pipeline still has limitations. Currently, since our SO-SMPL representation is based on SMPL-X model, our pipeline could not generate clothes types that do not follow SMPL-X’s topological structure, such as skirts and dresses. To extend to such clothes, we may follow [39] and utilize a parametric dress/skirt model as template instead of original SMPL-X. Another limitation is that our generated clothes do not follow any sewing patterns, and are without any physical attributes such as stretch, bending, and friction. Therefore, generating garments with sewing patterns and physical properties is a promising direction for exploration. We leave these potential tasks to future works.
@@ 7 Pipeline Details
7.1 Mask Templates
As mentioned in Sec. 4.1, we utilized SMPL-X[53] part segments to craft templates for 6 different types of clothes. The types of masks could be seen from Fig. 13.
@@ 7.1 Mask Templates
7.2 Negative Prompting
Simply using descriptive prompts can not guarantee the distinction between human body and clothes. For instance, even with descriptions like “unclothed” or “without accessories”, there could still be undesired extras on the human body. On the other hand for garment generation, the generated clothes sometimes contain undesired “human parts” like skin or even limbs. To address this problem, we added negative prompts for both body(e.g. “loose clothes, accessories”) and clothes generation(e.g. “skin, legs”), to encourage the predicted noise from Stable Diffusion to guide the model away from such entanglement. Specifically, the original classfier-free guidance[29] against unconditoned noise $\epsilon_{\phi}(\mathbf{I}_{t};t)$ is:
@@ 8 Additional Experiments
8.1 Experimental Details
All experiments for this paper are conducted on a single GeForce RTX 4090 GPU. For each human body mesh, it takes around 12GB GPU memory and 2.5 hours to generate. As for clothes, it takes around 19GB GPU memory and 2.5 hours of optimizing.
@@ 8.2 Avatar Gallery
A gallery of our generated avatars is presented in Fig. 14. Our method is able to generate highly detailed human body meshes.
@@ 8.3 Clothes Gallery
8.4 Additional Qualitative Comparisons with DreamWaltz [37]
We present comparisons with an additional state-of-the-art text-to-avatar method, namely DreamWaltz [37]. Visual comparisons are presented in Fig. 18 and Fig. 17.
As seen in Fig. 17, DreamWaltz [37] is capable of generating photorealistic plausible results, yet it suffers from over-smooth texture and noises due to its NeRF-based representation. Consequently in Fig. 18, noises move along with the joint motions, causing severe visual artifacts.
Figure 17: Comparisons of “A-pose” quality with DreamWaltz [37].
Figure 18: Comparisons of animation quality with DreamWaltz [37]. For a better visual comparison, please refer to our supplementary videos.

### v2

@@ 5.1 Generated Avatars & Clothes
Figure 5: A gallery of our generated clothed avatars. Our method can generate avatars with varied ethnicities, genders, and clothes.
As demonstrated in Fig. 5, our method generates diverse avatars and various types of clothes with detailed textures. Besides, since the clothes are generated on a base human body, they fit with the intended human naturally.
@@ 5.2 Quality Comparisons
We conduct comparisons on three aspects, static clothed-avatar quality, clothes quality, and animation quality. For the generation quality of clothed avatars, we compare with state-of-the-art text-to-3D avatar generation methods TEXture [76], AvatarCLIP [35] and TADA! [51]. As for clothes generation, we conduct comparisons with two of the best existing text-to-3D methods ProlificDreamer [94] and TextMesh [86]. We then compare animation results in the physical simulation environment [2] with the animatable meshes generated by TEXture [76] and TADA! [51].
@@
Figure 6: Quality comparison of the generated clothed avatars. We compare results from TEXture [76], AvatarCLIP [35], TADA! [51] and our method with the same prompts.
A visual comparison of the generated avatars is presented in Fig. 6. TEXTure [76] takes a fixed SMPL-X mesh as input and suffers from texture inconsistency under different views. AvatarCLIP [35] often produces geometry artifacts and poor facial details. TADA! [51] tends to generate unnatural geometric artifacts and unaligned outfits, e.g. a hull on the African-American woman’s belly and the white teenage girl’s yellow pants that contradicts the “beige legging” described by the prompt. This is because they use a simple text-to-image diffusion model and lack sufficient regularization on geometry. In comparison, using a pose-controlled diffusion model and proper regularization, our method can produce not only the most realistic human with detailed textures but also high-quality clothes that are semantically aligned with texts.
Clothes quality. We also evaluate the quality of the separated clothes. As demonstrated in Fig. 7, neither TextMesh [86] or ProlificDreamer [94] can generate meaningful clothes shapes or textures. Besides, they cannot control the position and the size of the generated clothes, and thus cannot be fitted to an avatar. In comparison, our method can produce clothes fitting to a specific avatar and also with high-quality textures and shapes.
Figure 7: Comparisons of clothes generation quality. Garments in each row are generated using the same prompts.
@@
Methods AvatarCLIP [35] TEXture [76] TADA [51] Ours
FID $\downarrow$ 166.0 132.8 90.2 84.9
CLIP-FID $\downarrow$ 34.9 42.8 27.4 26.2
CLIP Score $\uparrow$ 19.2 23.6 24.9 28.4
Table 1: Quantitative comparisons of generated “A”-pose avatar quality in FID, CLIP-FID, and CLIP score.
@@
Table 2: User study from 26 users in terms of visual quality, animation quality, and prompt alignment preference. The number means the percentage of users favoring a specific method.
5.3 Quantitative Comparisons
Accurate evaluation of text-to-3D generation is a challenging task, which is rarely discussed in existing works [98]. Here, we use 3 metrics to evaluate generation fidelity: FID [31, 67], CLIP-FID [48] and CLIP Score [30]. CLIP Score is the CLIP feature distance between the input prompts and the rendered images of the generated avatars. For FID and CLIP-FID, we measure the distance between the rendered images from our generated 3D avatars, and images generated from Stable Diffusion [77] using the same prompts. We evaluate all methods using 30 prompts and render 300 views for each prompt, resulting in 9000 images. As shown in Table 1, our method achieves the best results in all 3 metrics compared with previous methods, demonstrating the high visual quality and textual alignment of our pipeline.
5.4 User Study
We also conduct a comprehensive user study to evaluate both the generation quality and the animation capability of our pipeline. We design a questionnaire and present 16 full-body animations made by TEXture [76], TADA! [51], and our method. The survey is conducted among 26 researchers and artists from either academia or industry. The participants are asked to pick the one with the best semantic consistency, visual quality, and animation quality. Note that we ask participants to exclude animation quality when evaluating texture& geometry quality. As shown in Table  2, our model is the most preferred one in all aspects of generation quality.
@@ 5.5 Ablation Study
We first evaluate the effects of our albedo smoothness constraints and 2D-pose ControlNet guidance. In Fig. 9, we compare the rendered albedo and RGB images w/ or w/o our albedo smoothness constraint $\mathcal{L}_{a}$. As can be seen from the visual results, the albedo generated without $\mathcal{L}_{a}$ shows baked-in shadows and wrinkles, which leads to unsatisfactory visual artifacts in animations. In comparison, using $\mathcal{L}_{a}$ largely alleviates the problem in the albedo while still enabling the correct rendering of the shadows and wrinkles on RGB images.
In Fig. 9, we validate the effectiveness of OpenPose-based ControlNet as guidance during our human body generation. As can be observed from the results, generated avatars without ControlNet have over-smoothed textures and geometry. In contrast, our utilization of ControlNet produces significantly finer details in both geometry and texture.
Figure 9: Ablation study on our shading model (left) and ControlNet as guidance (right).
In Fig. 10 we demonstrate the necessity of using both guidances from clothes renders, $\{\mathbf{I}_{\text{c}},\mathbf{N}_{\text{c}}\}$, and clothed human renders, $\{\mathbf{I}_{\text{c+h}},\mathbf{N}_{\text{c+h}}\}$. Applying $\{\mathbf{I}_{\text{c+h}},\mathbf{N}_{\text{c+h}}\}$ without clothes leads to the incorrect generation of skin colors on the clothes meshes, as seen on the left of Fig. 10. On the other hand, with only clothes guidance, the garments are prone to learn texture and semantics incoherent with the human body, such as the back collar on the right of Fig. 10.
Figure 10: Ablation study on guidance from clothes renders$\{\mathbf{I}_{\text{c}},\mathbf{N}_{\text{c}}\}$ and clothed human renders $\{\mathbf{I}_{\text{c+h}},\mathbf{N}_{\text{c+h}}\}$.
@@ 5.6 Applications
Complex garments generation. Our SO-SMPL design can also be generalized to more complex garment types. For instance, by adding additional layers of offsets on top of the first clothing layer, we can generate more complex multi-layer garments, as shown in Fig. 11. Besides, leveraging BCNet [42], a parametric skirts model conditioned on SMPL parameters, our pipeline can be extended to generating clothes that do not follow SMPL-X’s topological structure such as skirts and dresses, as depicted in Fig. 11. More details can be found in the supplementary material.
Figure 11: Complex garments generation. Our pipeline could be generalized to generate more complex garments like skirts and multi-layered garments.
Virtual try-on. Our disentangled human-clothes representation inherently allows us to change the outfit of a certain avatar, or put the same clothes on different avatars. As depicted in Fig. 12, while previous methods like TADA! [51] also enable virtual try-on, they struggle to maintain consistency in clothing and human identity during the try-on process. As can be seen from the left half of Fig. 12, when changing clothes for the white teenage boy with TADA!, the facial details and body shape are changed; on the other hand, the colors and patterns of the clothes also changed when putting the same clothes on different avatars with TADA!. On the contrary, our disentangled representation supports fixing the human identity and clothes details when switching outfits.
@@ 6 Limitations&Conlusion
Limitations. Despite promising results, our pipeline still has limitations. Since images generated by diffusion models are of final shadings, it is difficult to get fully shadow-free material decomposition even with our shading model and albedo smoothness constraint. Besides, our results still suffer from cartoonish colors and over-smoothing textures, which is a common problem for distillation-based frameworks. Another limitation is the over-saturation of colors in our distillation-based framework. Although distillation techniques[94] have aimed to address this issue, it remains a challenge. Our offset-based representation also can not handle avatars with long hair, which requires advanced 3D representations[83]. What’s more, our generated clothes do not follow any sewing patterns and are without any physical attributes such as stretch, bending, and friction. Therefore, generating garments with sewing patterns is a promising direction for exploration. We leave these potential tasks to future works.
@@
Acknowledgements
We thank Tingting Liao, and Yukang Cao for their insightful suggestions. We also appreciate Ziyu Chen for fruitful discussions during the research process.
This work was partly supported by the Fundamental Research Funds for the Central Universities, 111 Project, China under Grants B07022 and (Sheitc) 150633, and Shanghai Key Laboratory of Digital Media Processing and Transmissions, China, and U.S. National Science Foundation CBET-2115405. This work is also partly supported by the Innovation and Technology Commission of the HKSAR Government under the ITSP-Platform grant (Ref: ITS/319/21FP) and the InnoHK initiative (TransGP project).
@@ 7 Pipeline Details
7.1 Garment Types
As mentioned in Sec. 4.1, we utilized SMPL-X[59] part segments to craft templates for 6 different types of clothes. The types of masks can be seen in Fig. 13. Note that our pipeline could be generalized to other shapes of garments by defining new masks, e.g. we could define a mask for feet areas to generate shoes.
@@ 7.1 Garment Types
As mentioned in the main paper, we could also leverage templates from BCNet [42] to generate skirts and short skirts. Specifically, we used the provided base template mesh along with its PCA-based linear displacement deformation. The PCA coefficients for the deformation could be produced through a two-layer MLP network conditioned on human body shape parameter $\beta$. As a result, a fitting skirt mesh template is acquired and could be used in our pipeline for offsets and texture optimization.
7.2 Prompting
Besides the input prompts, we also added extra descriptions after the prompts to enhance the performance. For human body generation, we added photorealistic, ultra-detailed, 8k uhd. For garments, we added wrinkleless smooth and flat. These descriptions are chosen empirically.
Simply using descriptive prompts can not guarantee the distinction between the human body and clothes. For instance, even with descriptions like “unclothed” or “without accessories”, there could still be undesired extras on the human body. On the other hand for garment generation, the generated clothes sometimes contain undesired “human parts” like skin or even limbs. To address this problem, we added negative prompts for both bodies (e.g. “loose clothes, accessories”) and clothes generation(e.g. “skin, legs”), to encourage the predicted noise from Stable Diffusion to guide the model away from such entanglement. Specifically, the original classifier-free guidance[33] against unconditioned noise $\epsilon_{\phi}(\mathbf{I}_{t};t)$ is:
@@ 8 Additional Experiments
8.1 Implementation Details
All experiments for this paper are conducted on a single GeForce RTX 4090 GPU. For each human body mesh, we optimize our model for 15000 steps, which takes around 12GB of GPU memory and 2.5 hours to generate. As for clothes, we optimize them for 12000 steps, which takes around 19GB of GPU memory and 2.5 hours of optimization. For human body optimization, we randomly sample around the center of the human body with an azimuth angle of [-180, 180] within a distance range of [1.25, 2.3], with a camera FoV of [45, 50]. We also add semantic close-ups on the face and hands region to enhance the fidelity of details. We adopt a similar strategy for clothes optimization, except that we introduce a position bias to coup with different garment types, e.g. move up the camera for upper garments. We set a learning rate of 0.0003 for offsets, 0.005 for texture, and 0.003 for body shape $\beta$. We also subdivide the original SMPL-X mesh to increase our model’s ability to represent geometric details, following TADA [51].
8.2 Evaluation Details
We use a digital form to conduct our user study. We show part of our user study form in Fig. 17. Three randomly ordered animation gifs made by different methods are played simultaneously. For each animation, three questions regarding text alignment, visual quality, and animation realism were asked. The criteria for the three metrics are described at the beginning of the form, in which we explicitly instruct participants to assess the visual quality of the avatars independently of the animation quality. Besides, we also conducted an additional user study among 18 users that compares our method with AvatarCLIP [35], in which our method is favored among all cases and categories. The avatars generated by AvatarCLIP exhibit noisy geometry and low-fidelity textures, as shown in Fig. 6.
Figure 17: Demonstration of our user study form.
For quantitative computations, we picked 30 prompts for generation, and we randomly sampled 300 camera views centered around the generated human body center. Then, we render all generated avatars with these same views, resulting in 9000 rendered images for each method. Close-up views on faces and hands are also sampled and paired with respective prompts, i.e. the face of …, the hand of …. We then used Stable Diffusion v1.5 [77] to generate 300 images for each prompt as the reference set. FID and Clean-FID between rendered images and the reference set are computed using the public tool provided by [67]. CLIP scores are measured directly between the input prompts and the rendered images using torchmetrics [64]. In tab.3, we also added additional prompts to make a more sufficient comparison between the most competitive concurrent work with our results.
8.3 Shape Editing
Our SMPL+offset-based model also allows us to edit the shape of generated characters and garments. To achieve this, we could simply modify the SMPL-X shape parameter $\beta$ of a character, and then add the garment offsets $\mathbf{O}_{\text{c}}$ on top of the edited character. As shown in Fig. 18, clothes generated by our methods could be changed according to different human shapes and sizes without re-generating. This flexibility helps us to manipulate the generated human body and allows generated clothes to be re-used on different avatars while maintaining fit with their body types.
Figure 18: Shape editing on clothes. SO-SMPL representation allows us to edit human shapes and fit clothes to them while maintaining the original patterns and details.
@@ 8.4 Avatar Gallery
A gallery of our generated avatars is presented in Fig. 14. Our method can generate highly detailed human body meshes.
@@ 8.5 Clothes Gallery
8.6 Additional Quantitative Comparisons
We added extra quantitative comparisons of “A”-pose avatar quality on our method and the most competitive concurent work TADA [51]. The experiment is conducted on 50 prompts with 300 renders each. As could be seen from Tab. 3, our method outperforms TADA [51] in all three metrics.
Methods TADA [51] Ours
FID $\downarrow$ 104.1 98.8
CLIP-FID $\downarrow$ 28.7 26.2
CLIP Score $\uparrow$ 24.7 27.8
Table 3: Additional comparisons of generated “A”-pose avatar quality in FID, CLIP-FID, and CLIP score.
8.7 Additional Qualitative Comparisons
We present comparisons with additional state-of-the-art animatable text-to-avatar methods, namely DreamWaltz [41] and HumanGaussian [55]. Visual comparisons are presented in Fig. 20 and Fig. 19.
As seen in Fig. 19, DreamWaltz [41] is capable of generating photorealistic plausible results, yet it suffers from over-smooth texture and noises due to its NeRF-based representation. Consequently in Fig. 20, noises move along with the joint motions, causing severe visual artifacts. A more recent work HumanGaussian [55] also suffers from geometric artifacts for its point-based Gaussian representation, resulting in noisy textures and motions.
Figure 19: Comparisons of “A-pose” quality with HumanGaussian [55] and DreamWaltz [41].
Figure 20: Comparisons of animation quality with HumanGaussian [55] and DreamWaltz [41]. For a better visual comparison, please refer to our supplementary videos.
9 Negative Impact
The advancement of text-driven avatar and clothing generation technologies carries with it potential risks such as privacy violations, intellectual property infringement, and the propagation of misinformation. Our human body generation method is particularly susceptible to malicious applications, which raises significant ethical concerns. We emphasize the necessity for transparency in further development and applications of this technology. In addition, we advocate for the establishment of robust oversight measures to ensure that the utilization of these advancements does not harm individuals or society.